# AI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions

### v1

@@ Abstract
This paper provides a comprehensive analysis of cognitive biases in forensics and digital forensics, examining their implications for decision-making processes in these fields. It explores the various types of cognitive biases that may arise during forensic investigations and digital forensic analyses, such as confirmation bias, expectation bias, overconfidence in errors, contextual bias, and attributional biases. It also evaluates existing methods and techniques used to mitigate cognitive biases in these contexts, assessing the effectiveness of interventions aimed at reducing biases and improving decision-making outcomes. Additionally, this paper introduces a new cognitive bias, called “impostor bias”, that may affect the use of generative Artificial Intelligence (AI) tools in forensics and digital forensics. The impostor bias is the tendency to doubt the authenticity or validity of the output generated by AI tools, such as deepfakes, in the form of audio, images, and videos. This bias may lead to erroneous judgments or false accusations, undermining the reliability and credibility of forensic evidence. The paper discusses the potential causes and consequences of the impostor bias, and suggests some strategies to prevent or counteract it. By addressing these topics, this paper seeks to offer valuable insights into understanding cognitive biases in forensic practices and provide recommendations for future research and practical applications to enhance the objectivity and validity of forensic investigations.
@@ 1 Introduction
In the realm of forensic sciences, perhaps more than in other disciplines, the so-called objectivity of judgment in analyzing data and interpreting them for justice purposes is one of the crucial aspects. Beyond the specific technical aspects of individual disciplines, the authoritativeness of studies, and the competence of experts, there is a need to be informed about the so-called cognitive biases. Cognitive biases, rather than being deficits, are systematic preferences that influence the way we process, select, and retain information (Grisham et al., 2014; Lester et al., 2011). These biases can significantly impact our judgment and decision-making processes.
The tendency to jump to conclusions involves making hasty judgments or drawing premature conclusions without adequate evidence (Warman et al., 2013). Confirmation bias is another common cognitive bias where individuals selectively pay attention to information that aligns with their pre-existing beliefs, often disregarding contradictory evidence (Peters, 2020). Expectation bias, also known as experimenter’s bias, occurs when an individual’s anticipated outcome influences the actual outcome observed (Forensic Science Regulator, 2020). Overconfidence in errors is a related bias where individuals exhibit unwarranted certainty in their judgments and predictions, even when proven incorrect (Han, 2020; Sauvé et al., 2020). Contextual bias refers to situations where an individual’s judgment is influenced, either consciously or subconsciously, by additional information beyond what is being directly evaluated (Forensic Science Regulator, 2020). Attributional biases involve the unfair attribution of external events or experiences to oneself or others, thereby influencing our perception of causality (Sauvé et al., 2020). Cognitive biases can have both positive and negative effects, depending on the context and situation. They can facilitate swift decision-making when time is critical but can also lead to poor decisions and adverse outcomes (Berthet, 2022). Cognitive biases have been found to significantly influence decision-making processes in various professional fields such as management, finance, medicine, law, and psychology (Acciarini et al., 2021; Berthet, 2022; Neal and Grisso, 2014). Overconfidence bias, in particular, is a recurring bias that impacts decisions in these areas (Grežo, 2021; Wu et al., 2012). A summary of these biases can be found in Table 1.
Bias Description
Jump to Conclusions
Making quick judgments or drawing premature conclusions without sufficient evidence.
Confirmation Bias
Selectively paying attention to information that confirms our pre-existing beliefs while ignoring contradictory evidence.
Expectation Bias
An individual’s anticipated outcome influences the actual outcome that is observed.
Overconfidence in Errors
Individuals display an unwarranted certainty in their judgments and predictions, even when they are proven wrong.
Contextual Bias
An individual’s judgment is influenced (either consciously or subconsciously) by additional information beyond what is being directly evaluated.
Attributional Biases
Unfairly attributing external events or experiences to oneself or others, influencing our perception of causality.
Table 1: Summary of Cognitive Biases.
This paper delves into the influence of cognitive biases in the realm of forensics and digital forensics, with a particular focus on the detection of deepfakes. Deepfakes, which are Artificial Intelligence (AI)-generated images, videos, and audio, have a substantial impact on these fields. They can shape attitudes even when viewers are cognizant of the fact that they are viewing deepfakes, a phenomenon attributable to cognitive biases. Deepfakes pose threats such as the manipulation of public opinion and the impersonation of individuals, which could potentially lead to financial and reputational damage. In this context, we will introduce the concept of the impostor bias, which we define as an inherent distrust concerning the authenticity of multimedia elements like videos, photos, and audio. This distrust stems from the awareness that these elements can be realistically generated by AI models. Effective detection of deepfake products is crucial to prevent falling prey to the impostor bias and to avoid the potential confusion between real and fake multimedia products. Such confusion can lead to erroneous decisions and perceptions across all layers of multimedia evaluation in digital forensics. The challenges presented by deepfakes necessitate ongoing research and the development of advanced solutions. Consequently, we will analyze the most recent and efficient deepfake detection systems to aid operators and investigators in distinguishing between real and fake multimedia.
@@ 2 Cognitive Biases Impact on Forensic Sciences
Cognitive biases have been a concern in forensic science since 1984, when Larry Miller put forth a paper discussing the presence of bias in forensic document examiners (Stoel et al., 2014). He suggested the introduction of procedural modifications to reduce cognitive bias, which could potentially result in incorrect outcomes. These biases can significantly impact expert judgments and the criminal justice process (Stoel et al., 2014; Dror and Rosenthal, 2008; Neal and Grisso, 2014), are influenced by various factors, and can lead to errors and misinterpretation of evidence (Kassin et al., 2013; Cooper and Meterko, 2019; Bhadra, 2021). Both forensic experts and law enforcement professionals are susceptible to these biases, including confirmation bias (Thompson and Newman, 2015; Meterko and Cooper, 2022). The sources of bias can be categorized into three groups related to the case, the analyst, and human nature (Dror, 2020). Seven primary factors can introduce bias within the laboratory setting (Jeanguenat et al., 2017) (As synthetically sketched in Figure 1).
Figure 1: Seven potential sources of bias that could impact forensic decision-making.
Gardner et al. (2019) found that task-irrelevant information can bias forensic analysts’ decisions. Similarly, Nakhaeizadeh et al. (2014) discovered that irrelevant information can lead to confirmation bias in forensic anthropology. Dror et al. (2021) found that base-rate neglect could bias forensic pathologists’ decisions in child death cases. In DNA forensics, Jeanguenat et al. (2017) found that suspect-driven bias and rare alleles can influence interpretation. Neal et al. (2022) conducted a systematic review on cognitive biases and debiasing techniques in forensic mental health, finding significant bias effects. Lastly, Stevenage and Bennett (2017) found that irrelevant DNA test outcomes could bias fingerprint matching tasks, confirming the presence of contextual bias.
@@ 3 Exploring Cognitive Biases in Digital Forensics
The definition of Digital Forensic Science often referred to is the one from the Digital Forensic Research Workshop (DFRWS) in Palmer et al. (2001): “The use of scientifically derived and proven methods toward the preservation, collection, validation, identification, analysis, interpretation, documentation and presentation of digital evidence derived from digital sources for the purpose of facilitating or furthering the reconstruction of events found to be criminal, or helping to anticipate unauthorized actions shown to be disruptive to planned operations”. A key area within this field involves the proper acquisition of digital content such as images, videos, and audio to produce evidence for forensic investigations. Multimedia forensics focuses on verifying the authenticity of data and reconstructing the history of an image since its acquisition (Battiato et al., 2016; Fabio Arceri et al., 2023; Giudice et al., 2017; Piva, 2013). In the realm of digital forensics, cognitive biases can significantly impact the interpretation of data.
@@ 3.1 Confirmation Bias in Text and Face Recognition
Fontani’s blog post on Amped Software depicted a hypothetical situation featuring two characters, John and Lucy (Figure 2(Fontani, 2021). In this scenario, John inadvertently influenced Lucy’s interpretation of a license plate by prematurely sharing his own interpretation. This phenomenon, known as confirmation bias, may lead Lucy to unconsciously process pixels and select frames that align with John’s interpretation. The post further delved into the intricacies of face comparison, noting that varying processing techniques can result in significantly different facial appearances. It underscored the necessity of withholding the suspect’s face from the examiner prior to the enhancement process to prevent unconscious bias towards a match. Fontani also emphasized that different processing techniques can significantly alter facial appearances during enhancement (Figure 3). Therefore, it’s crucial that the examiner doesn’t see the suspect’s face before the enhancement process to avoid unconsciously adjusting the enhancement to create a match. This unconscious adjustment is a form of cognitive bias.
@@ 3.2 Image Processing Could Lead to Pareidolia
The enhancement of images extracted from video cameras can lead to a degradation of the overall quality of information. This is due to factors such as excessive compression, distance from the recording plane, and limited overall resolution. International best practices suggest verifying on a case-by-case basis whether the level of information is sufficient to extract useful data for investigations (Khin et al., 2020). However, when examining low-contrast images that present pseudo-random visual patterns after an initial enhancement process, it is crucial to mitigate the risk of pareidolia. This bias is particularly potent when the object of interest refers to “human faces” or more generally to “letters/numbers” or known human structures (Wang and Yang, 2018; Zhou and Meng, 2020a). Pareidolia is a subconscious illusion that tends to associate random shapes with known forms, especially human figures and faces. Classic examples include seeing animals or human faces in clouds, or a human face on the moon. In forensic investigations, we also suggest to entrust the analysis and interpretation to automatic methods or experts who can follow a “blind testing” approach, i.e., an interpretation detached from the knowledge of details and the reference context. This helps to ensure the search for “certain” evidence is as rigorous and unbiased as possible.
@@
Zhou and Meng (2020b) discussed how some individuals exhibit a looser decision criterion for detecting faces, making them more prone to perceive faces where none exist. This relates to a concept in signal detection theory known as response bias (Nguyen and Beins, 2013). In digital forensics, examiners may fall prey to response biases when analyzing ambiguous digital evidence, predisposing them to validate or dismiss forensic hypotheses based on non-diagnostic features. Just as some are more likely to see faces in random patterns due to biases in how they set thresholds for face judgments, forensic analysts could have biases influencing how strictly they apply standards of evidence to digital artifacts. Understanding individual differences in cognitive biases like threshold placement could help address potential sources of error and increase objectivity in digital forensic examinations.
Furthermore, we present a case study which involved the analysis of surveillance camera footage to determine the presence of a passenger in a vehicle involved in a legal case. The analysis included scrutiny of various cameras and raised concerns about the validity of conclusions drawn by experts. The frames are part of a detailed analysis of the vehicle’s passages from different surveillance cameras, considering image overlaps and real passage times, including sunset. The aim was to evaluate the methodologies and conclusions of technical consultants and expert witnesses regarding the analysis of vehicle images, highlighting the lack of scientific rigor and proposing new analyses suggesting the absence of a passenger in the vehicle. Specifically, Figure 6 shows a sequence of frames that have been processed by a technical consultant. The frames demonstrate a clear reflection effect that disappears as the vehicle moves. The effect is attributed to the ease with which contour lines can introduce aberrations. The frames also exhibit strong chromatic variability due to video compression, which is known to heavily penalize the color component.
@@
The technical advisor claimed to see the image of human faces in such images, a factor induced by confirmation bias: in reality, what is present in the images are pseudo-random blobs and their temporal hold is instead to be traced to a simpler and more obvious reflection. To demonstrate this, we conducted an experiment with college students. In the experiment, students were presented with frames from surveillance camera footage and tasked with identifying and interpreting the images. Crucially, no prior information was provided to the students to prevent any preconceived notions or confirmation biases from influencing their observations. The primary focus was on the students’ ability to discern whether a vehicle contained a passenger or not. The experiment was designed to rigorously assess the students’ cognitive perception, comprehension of the images, and their capacity to make accurate identifications solely based on the visual data at hand. This approach underscores the importance of maintaining an unbiased perspective in observational tasks. Table 2 summarizes the outcome of this analysis process, which reports a number of responses of 165 (15 subjects multiplied by the number of images viewed) in which in no case was the ”certain” presence of a human face evidenced.
@@
Those who answered ”other” identified:
@@
Those who answered ”other” identified:
@@
Those who answered ”other” identified:
@@
Those who answered ”other” identified:
@@
Those who answered ”other” identified:
@@
Those who answered ”other” identified:
@@
Table 2: Table representing students’ evaluation of the various images cutted from Figure 6.
4 Bias Mitigating Strategies
Mitigating these cognitive biases could significantly enhance decision-making across society, promoting long-term human well-being. Extensive research has been conducted on whether and how these biases can be mitigated (Acciarini et al., 2021).
Korteling et al. (2021) conducted a systematic review of the literature on the retention and transfer of bias mitigation interventions, which yielded 12 peer-reviewed studies that adequately studied retention over a period of at least 14 days or transfer to different tasks and contexts. Most of these studies indicated that game-based interventions were effective after the retention interval and were more effective than video interventions. One study found indications of transfer of bias mitigation training across contexts.
4.1 Biases Mitigating Strategies in Forensics

### v2

# GenAI Mirage: The Impostor Bias and the Deepfake Detection Challenge in the Era of Artificial Illusions

@@ Abstract
This paper examines the impact of cognitive biases on decision-making in forensics and digital forensics, exploring biases such as confirmation bias, anchoring bias, and hindsight bias. It assesses existing methods to mitigate biases and improve decision-making, introducing the novel “Impostor Bias”, which arises as a systematic tendency to question the authenticity of multimedia content, such as audio, images, and videos, often assuming they are generated by AI tools. This bias goes beyond evaluators’ knowledge levels, as it can lead to erroneous judgments and false accusations, undermining the reliability and credibility of forensic evidence. Impostor Bias stems from an a priori assumption rather than an objective content assessment, and its impact is expected to grow with the increasing realism of AI-generated multimedia products. The paper discusses the potential causes and consequences of Impostor Bias, suggesting strategies for prevention and counteraction. By addressing these topics, this paper aims to provide valuable insights, enhance the objectivity and validity of forensic investigations, and offer recommendations for future research and practical applications to ensure the integrity and reliability of forensic practices.
@@
journal: Forensic Science International: Digital Investigation
@@ 1 Introduction
In forensic sciences, the objectivity of judgment in analyzing data for justice purposes is paramount. Beyond technical expertise, awareness of cognitive biases is crucial. These biases, rather than being deficits, are systematic preferences that influence the way we process, select, and retain information (Lester et al., 2011; Grisham et al., 2014). They can have both positive and negative effects, depending on the context and situation, facilitating swift decision-making when time is critical but also leading to poor decisions and adverse outcomes (Meterko and Cooper, 2022; Berthet, 2022).
Some of these biases have been identified as a significant factor impacting the objectivity and accuracy of forensic science (Bhadra, 2021). For example, law enforcement professionals showed vulnerable to confirmation bias (Meterko and Cooper, 2022), which is a tendency to search for, interpret, and remember information in a way that confirms one’s preexisting beliefs or hypotheses (Nickerson, 1998). Another recurring bias in forensics is the anchoring bias or effect (Edmond et al., 2015), which occurs when an individual relies too heavily on an initial piece of information (the “anchor”) when making decisions (Chapman and Johnson, 1994, 1999). One more relevant bias related to forensic science is the hindsight bias (Giroux et al., 2016), which occurs when people believe that an event is more predictable after it becomes known, involving memory distortion, beliefs about objective likelihoods, and subjective beliefs about one’s own prediction abilities (Roese and Vohs, 2012). The discipline of digital forensics is not exempt from these challenges (Sunde and Dror, 2019). Consequently, an increasing number of scholarly investigations are focusing on this particular field (Sunde and Dror, 2021).
This paper explores the impact of cognitive biases in forensics and digital forensics, with a focus on deepfakes and Artificial Intelligence (AI)-generated multimedia content, which pose threats such as manipulating public opinion and impersonating individuals. We introduce the Impostor Bias, an inherent distrust of multimedia authenticity due to the prevalence of AI-generated content. Effective detection of deepfakes is crucial to prevent confusion between real and fake multimedia, which can lead to erroneous decisions. We analyze advanced deepfake detection systems to aid operators in distinguishing authentic content, addressing the challenges posed by deepfakes, and ensuring accurate multimedia evaluation in digital forensics.
@@ 2 Cognitive Biases Impact on Forensic Sciences
Cognitive biases have been a concern in forensic science since 1984, when Larry Miller published a work discussing the presence of bias in forensic document examiners (Stoel et al., 2014). He suggested the introduction of procedural modifications to reduce cognitive bias, which could potentially result in incorrect outcomes.
These biases can significantly impact expert judgments and the criminal justice process (Stoel et al., 2014; Dror and Rosenthal, 2008; Neal and Grisso, 2014), are influenced by various factors, and can lead to errors and misinterpretation of evidence (Kassin et al., 2013; Cooper and Meterko, 2019; Bhadra, 2021). Both forensic experts and law enforcement professionals are susceptible to these biases, and one of the most common of them is the confirmation bias (Moser, 2013; Thompson and Newman, 2015; van den Eeden et al., 2019; Cooper and Meterko, 2019; Meterko and Cooper, 2022): if a forensic analyst believes a suspect is guilty, they might unconsciously interpret ambiguous evidence as incriminating (Kassin et al., 2013; van den Eeden et al., 2019)Gardner et al. (2019) found that task-irrelevant information can bias forensic analysts’ decisions, suggesting that extraneous details can inadvertently sway the interpretation of evidence. This is echoed by  Nakhaeizadeh et al. (2014), who discovered that irrelevant information can lead to confirmation bias in forensic anthropology, potentially leading to skewed conclusions based on pre-existing beliefs rather than objective evidence. Dror et al. (2021) further explored this issue, finding that base-rate neglect could bias forensic pathologists’ decisions in child death cases. This suggests that statistical information about the prevalence of certain causes of death may be overlooked, leading to potential misinterpretations. In the realm of DNA forensics, Jeanguenat et al. (2017) found that suspect-driven bias and the presence of rare alleles can influence interpretation. This highlights the potential for preconceived notions about a suspect, as well as the rarity of certain genetic markers, to affect the analysis of DNA evidence. In a study by Douglass et al. (2023), evaluators successfully differentiated accurate from inaccurate witnesses based on videos of identification procedures alone, but their ability to discern accuracy was disrupted when extraneous incriminating evidence was also provided. This aligns with confirmation bias, where evaluators tend to favor information that confirms their existing beliefs or expectations, even when it contradicts objective evidence. Regarding the anchoring bias, in a forensic context, an analyst might give undue weight to the first piece of evidence they examine, which could skew their interpretation of subsequent evidence (Edmond et al., 2015; Meterko and Cooper, 2022). For the hindsight bias, this could lead to overconfidence in the accuracy of a forensic analysis after a suspect has been identified (Giroux et al., 2016; Beltrani et al., 2018; Meterko and Cooper, 2022)Neal et al. (2022) conducted a systematic review on cognitive biases and debiasing techniques in forensic mental health, finding significant bias effects specifically for confirmation and hindsight bias. Lastly, Stevenage and Bennett (2017) found that irrelevant DNA test outcomes could bias fingerprint matching tasks, confirming the presence of contextual bias. This suggests that unrelated test results can influence the interpretation of fingerprint evidence.
The sources of bias can be categorized into three groups related to the case, the analyst, and human nature. Inspired by the work of Dror et al. (Dror, 2020), we define a taxonomy of potential sources of bias. These biases can introduce cognitive distortions into the processes of sampling, observing, strategizing tests, analyzing, and drawing conclusions, even when conducted by experts. This taxonomy is synthetically sketched in Figure 1.
Figure 1: Eight potential sources of bias that could impact forensic decision-making.
@@ 3 Exploring Cognitive Biases in Digital Forensics
The definition of Digital Forensic Science often referred to is the one from the Digital Forensic Research Workshop (DFRWS) in Palmer et al. (2001): “The use of scientifically derived and proven methods toward the preservation, collection, validation, identification, analysis, interpretation, documentation and presentation of digital evidence derived from digital sources for the purpose of facilitating or furthering the reconstruction of events found to be criminal, or helping to anticipate unauthorized actions shown to be disruptive to planned operations”. A key area within this field involves the proper acquisition of digital content such as images, videos, and audio to produce evidence for forensic investigations. Multimedia forensics focuses on verifying the authenticity of data and reconstructing the history of an image since its acquisition (Battiato et al., 2016; Arceri et al., 2023; Giudice et al., 2017; Piva, 2013).
In the realm of digital forensics, cognitive bias emerges as a subtle yet potent force that can shape the outcomes of investigations. Sunde and Dror (2021) explored the susceptibility of digital forensics examiners to bias, revealing how preconceived notions and contextual information can skew their observations and interpretations. Despite the digital evidence’s facade of objectivity, the human factor introduces variability, leading to inconsistent conclusions among experts analyzing identical datasets (Sunde and Dror, 2021).
@@ 3.1 Confirmation Bias in Text and Face Recognition
Fontani’s blog post on Amped Software depicted a hypothetical situation featuring two characters, John and Lucy (Figure 2(Fontani, 2021). In this scenario, John inadvertently influenced Lucy’s interpretation of a license plate by prematurely sharing his own interpretation. The cognitive bias present in this example could be the confirmation bias, since this latter is a type of cognitive bias where individuals are more likely to seek out, interpret, and remember information that confirms their pre-existing belief (Nickerson, 1998; Cooper and Meterko, 2019; Meterko and Cooper, 2022). John’s premature sharing of his interpretation could lead Lucy to interpret the license plate in the same way, confirming John’s interpretation rather than considering other possible interpretations. It may lead Lucy to unconsciously process pixels and select frames that align with John’s interpretation. The post further exploited into the intricacies of face comparison, noting that varying processing techniques can result in significantly different facial appearances. It underscored the necessity of withholding the suspect’s face from the examiner prior to the enhancement process to prevent unconscious bias towards a match. Fontani also emphasized that different processing techniques can significantly alter facial appearances during enhancement (Figure 3). Therefore, it’s crucial that the examiner doesn’t see the suspect’s face before the enhancement process to avoid unconsciously adjusting the enhancement to create a match.
@@ 3.2 Image Processing Could Lead to Pareidolia
The enhancement of images extracted from video cameras can indeed lead to a degradation of the overall quality of information. This degradation can be attributed to factors such as excessive compression (Maity et al., 2023), distance from the recording plane (Wang et al., 2023), and limited overall resolution (Maity et al., 2023; Wang et al., 2023). International best practices suggest verifying on a case-by-case basis whether the level of information is sufficient to extract useful data for investigations (Tenopir et al., 2020; Soltani and Nikou, 2020).
@@
However, when examining low-contrast images that present pseudo-random visual patterns after an initial enhancement process, it is crucial to mitigate the risk of pareidolia. This bias is particularly potent when the object of interest refers to “human faces” or more generally to “letters/numbers” or known human structures (Wang and Yang, 2018; Zhou and Meng, 2020). Pareidolia is a subconscious illusion that tends to associate random shapes with known forms, especially human figures and faces. Classic examples include seeing animals or human faces in clouds, or a human face on the moon. In forensic investigations, we also suggest to entrust the analysis and interpretation to automatic methods (Solanke and Biasiotti, 2022) or experts who can follow a “blind testing” approach, i.e., an interpretation detached from the knowledge of details and the reference context (Cowan and Koppl, 2011; Servick, 2015). This helps to ensure the search for “certain” evidence is as rigorous and unbiased as possible.
@@
Zhou and Meng (2020) discussed how some individuals exhibit a looser decision criterion for detecting faces, making them more prone to perceive faces where none exist. This relates to a concept in signal detection theory known as response bias (Nguyen and Beins, 2013). In digital forensics, examiners may fall prey to response biases when analyzing ambiguous digital evidence, predisposing them to validate or dismiss forensic hypotheses based on non-diagnostic features. Just as some are more likely to see faces in random patterns due to biases in how they set thresholds for face judgments, forensic analysts could have biases influencing how strictly they apply standards of evidence to digital artifacts. Understanding individual differences in cognitive biases like threshold placement could help address potential sources of error and increase objectivity in digital forensic examinations (Berthet, 2021; Horsman, 2024).
3.3 Case Study: Confirmation Bias and Pareidolia in Surveillance Camera Footage
We present a case study with the objective of determining the presence of a passenger in a vehicle involved in a murder criminal case, through the analysis of surveillance camera footage. The methodology included the scrutiny of various cameras and a detailed analysis of the vehicle’s passages from different surveillance cameras, considering image overlaps and real passage times, including sunset. The data for this study is primarily derived from the surveillance camera footage. Figure 6 shows the only sequence of frames (among all the collected surveillance footage) in which the passenger’s presence is doubtful, and that has been processed by a technical consultant. These frames demonstrate a clear reflection effect that disappears as the vehicle moves, and exhibit strong chromatic variability due to video compression. The technical advisor claimed to discern a human face in these images, a perception potentially induced by confirmation bias, as his intervention was required precisely to identify the possible presence of the passenger. Specifically, on the passage of the car identified by the technical advisor he identifies with certainty, in the first frames of the passage, a face as a “clear” silhouette traceable to a passenger placed in the right front seat. A clear spot can be seen that he traces back to a silhouette. This spot disappears completely in the second part of the passage and then reappears but in inverted colors, that is, it becomes dark. In reality, the images contain pseudo-random blobs, and their temporal persistence can be attributed to a simpler and more evident reflection. The study highlights the lack of scientific rigor in these approaches and proposes new analyses that suggest the absence of a passenger in the vehicle.
@@ 3.3 Case Study: Confirmation Bias and Pareidolia in Surveillance Camera Footage
Those who answered “other” identified:
@@
Those who answered “other” identified:
@@
Those who answered “other” identified:
@@
Those who answered “other” identified:
@@
Those who answered “other” identified:
@@
Those who answered “other” identified:
@@
Table 1: Table representing students’ evaluation of the various images extracted from Figure 6.
An experiment was conducted involving college students. The students were anonymously and without context presented with carefully selected frames from the surveillance camera footage, focusing on the most controversial and doubtful portions. The images shown to the students were a representative subset of the entire set of images extracted and processed by the technical advisor (Figure6). To ensure impartiality in their observations, no preliminary information was provided to the students. As part of the experiment, a questionnaire was administered for each image, consisting of the question: “What do you observe in this image?”, with response options such as “Nothing” and “Other,” along with an open text field for further clarification. This design encouraged a spectrum of responses and allowed students to give concise yet descriptive explanations. A total of 15 subjects participated in evaluating the selected images, resulting in 165 responses, as summarized in Table 1. Notably, only two of the responses conveyed a ”certain” identification of a human face. This underscores the necessity of meticulous and unbiased analysis when interpreting surveillance footage. The primary objective of the experiment was to assess the students’ ability to detect the presence of a passenger in the vehicle. By presenting them with ambiguous visuals, the experiment aimed to understand their perceptual limitations and biases. The absence of explicit certainty measurement in the questionnaire was addressed by considering the students’ chosen responses as an indication of their certainty levels. This highlights the importance of rigorous and unbiased analysis in interpreting surveillance footage. The study emphasizes the need to maintain an objective perspective in observational tasks (Altmann, 1974) and underscores the value of scientific rigor in such analyses.
4 Strategies for Mitigating Cognitive Bias
Among other fields, software engineering is currently experiencing a significant gap in the area of cognitive bias mitigation techniques, with a notable lack of both practical strategies and theoretical foundations (Mohanani et al., 2017). However, other fields have seen success in this area. For instance, in social work, the use of a nomogram tool and an online training course has been shown to effectively mitigate cognitive bias, leading to improvements in the accuracy of clinical reasoning (Featherston et al., 2019). Despite these advancements, it’s important to note that there is currently insufficient evidence to suggest that cognitive bias mitigation interventions significantly improve decision-making in real-life situations (Korteling et al., 2021).
Nevertheless, the potential benefits of countering cognitive biases are clear. In healthcare, for example, addressing cognitive biases in decision-making can help reduce low-value care and enhance the impact of campaigns aimed at reducing such care (Scott et al., 2017). In the realm of gaming, the MACBETH serious game has been found to effectively mitigate cognitive biases such as the fundamental attribution error (Miller and Lawson, 1989) and confirmation bias. The game’s effectiveness is further enhanced through explicit instruction and repetitive play, which serve to reinforce learning (Dunbar et al., 2014).
4.1 Strategies for Mitigating Cognitive Bias in Forensic Science